OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box, in slope-intercept form, without skipping any spaces. m = 3, point (2, 1)

OpenStudy (anonymous):

Slope-intercept form: y = mx + b m = slope b = y-intercept y = 3x + b Solve for b by plugging in (2, 1) 1 = 3(2) + b 1 = 6 + b b = -5 y = 3x - 5

OpenStudy (anonymous):

Thank you!